# Sublogarithmic Distributed Vertex Coloring with Optimal Number of Colors

### 30 Mar 2026 v1

@@ 5 Coloring With Much Slack
The goal of this section is to prove Lemma 4.5, that is, we aim to color an induced subgraph $H$ of $F-\bigcup_{i}A_{i}$ that is $\Pi$-ous while at the same time maintaining CC for all uncolored cliques. Recall that $H$ is $\Pi$-ous if each node has sufficient slack in $H$ compared to a given upper bound $U\geqslant\Delta^{1/4}$ on the external degree of uncolored cliques. The coloring process consists of two primary components. The first one is a random color trial for all uncolored vertices, applied for $O(\log\Delta)$ iterations, where each iteration lowers the uncolored degree of each vertex by a constant factor in expectation. The properties of the resulting partial coloring after a single color trial are summarized in Lemma 5.1 below. The second component summarized in Lemma 5.2 below is then a multi color trial that colors all remaining vertices of the graph instance to be colored.
@@
    CC is maintained for all uncolored cliques.
@@
Let $H$ be an uncolored induced subgraph in which every vertex has $\Delta^{9/20}$ slack and each uncolored vertex in each $A_{i}$ has at most $\Delta^{1/10}$ neighbors in $H$. There is a $\widetilde{O}(\log^{3}\log n)$-round $\mathsf{LOCAL}$ algorithm that, w.h.p., colors $H$ while maintaining CC for every uncolored clique $A_{i}$. When $\Delta\geqslant(\log n)^{50}$, the algorithm runs in $O(1)$ rounds.
@@
For sufficiently large constant $c$, this is at most $\Delta^{1/10}$. Now let $H^{\prime}$ be the subgraph of the initial graph $H$ that contains all the yet uncolored vertices after $T$ iterations of Lemma 5.1. Then the graph $H^{\prime}$ fulfills the preconditions of Lemma 5.2, which require that every vertex has $\Delta^{9/20}$ slack and each uncolored vertex in each $A_{i}$ has at most $\Delta^{1/10}$ neighbors in the remaining uncolored graph. The application of Lemma 5.2 to $H^{\prime}$ completes the coloring of $H$. The CC is preserved for all uncolored cliques as it is guaranteed by the outcome of Lemmas 5.1 and 5.2. The runtime consists of the $O(\log\Delta)$ iterations from Lemma 5.1 that take $\widetilde{O}(\log^{3}\log n)$ rounds each and $\widetilde{O}(\log^{3}\log n)$ rounds for the one application of Lemma 5.2. This completes the proof.
5.1 Iterated Random Color Trial with CC (Lemma 5.1)
The goal of this section is to prove Lemma 5.1. We present a $\operatorname{\text{{\rm poly}}}(\log\log n)$ algorithm that partially colors the graph $H$ while maintaining CC and additionally reduces the uncolored degree of vertices in $H$ and $\bigcup_{i}A_{i}$ below $\Delta^{1/10}$.
The analysis focuses on a single iteration of Random Color Trial (Rct). We show that in each iteration of the Rct procedure the uncolored degree of a vertex drops by a factor of (11/180)(1-1/180) compared to the previous iteration while maintaining CC. We set up an LLL that, if solved correctly, yields a partial coloring guaranteeing this drop in the uncolored degree. The random process underlying our LLL is described in Algorithm 2; we call the color $\psi(v)$ the candidate color of $v$.
@@ 2 Each activated vertex $v$ independently samples a uniform available color $\psi(v)$
$E_{CC}(i)$: CC is not maintained for some color w.r.t. $A_{i}$.
@@
Consider any uncolored clique $A_{i}$ and any color $x$. The relevant collection of subsets is formed by the external neighborhoods of each vertex $v\in A_{i}$. By Lemma 4.1-(a), we have $|A_{i}|<\Delta$. We ignore vertices of $\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ for their contribution is not accounted for in CC. Hence, no relevant vertex of $H$ belongs to more than $2\Delta^{9/10}$ such subsets. Since $H^{\prime}$ is $\Pi$-ous, each vertex $v$ of $H^{\prime}$ picks color $x$ as candidate color w.p. at most $1/(\deg_{H^{\prime}}(v)+U\cdot\Delta^{0.22}))\leqslant 1/(Q\cdot\Delta^{1/5})$ for $Q=U$. As candidate color choices are independent for different vertices, the probability that $\ell$ vertices pick color $x$ is at most $1/(Q\cdot\Delta^{1/5})^{\ell}$. Therefore, by Lemma 4.4, CC fails for clique $A_{i}$ and color $x$ w.p. at most $\exp(-\Delta^{1/40})$. The lemma follows by the union bound over a total of at most $\Delta$ different colors.
@@
As detailed in CC, there is a fresh budget of $CC(i,x)<\Delta^{\frac{37}{40}}$ for each color $x$, each clique $A_{i}$, and also for each coloring step of the algorithm. In particular, in Algorithms 3 and 3 of Algorithm 3 the events $E_{CC}(i)$ get a separate budget of $\Delta^{\frac{37}{40}}$ each.
@@
Step (S1) holds with $c_{1}=5$ due to 5.3 (and because $F$ includes the cliques): a vertex $v$ holds variables used only by events $E(u)$ with $u\in N_{H}^{\leqslant 2}(v)$ (which has diameter at most 4) or events $E_{CC}(i)$ with $A_{i}$ with $\operatorname{dist}_{G}(v,A_{i})\leqslant 2$ (which has diameter at most 5 in $F$). Step (S2) requires a bound $\Delta(F)^{c_{2}}$ for some constant $c_{2}$ on the number of events in which each variable appears and it holds with $c_{2}=2$ as each variable is only used by events with variables in its $2$-hop neighborhood. Step (S3) requires that for each event $A\in\mathcal{A}$ we can determine $1(A\in\mathcal{A}^{\prime})$ by only evaluating the variables in $\operatorname{vbl}(A)$ which immediately holds by the definition of $\mathcal{A}^{\prime}$. Step (S4) requires an upper bound of $1/\Delta(F)^{c_{4}}$ on the probability for each bad event $A\in\mathcal{A}$ to be contained in $\mathcal{A}^{\prime}$ where $c_{1}$ is an arbitrary constant satisfying $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1$. By Lemmas 5.4 and 5.5 this probability is at most $\Delta(G)\cdot\exp(-\Delta(G)^{1/40})$. As $\Delta(G)=\Theta(\Delta(F))$ Step (S4) holds for $c_{4}$ (as chosen above) as long as $\Delta(G)$ is larger than a sufficiently large absolute constant $\Delta_{0}$. The set of events included in the LLL of Algorithm 3 are all included in $\mathcal{B}^{\prime}$ (with $c_{\mathcal{B}^{\prime}}$ defined above) as vertices with distance at most $2$ to $\operatorname{vbl}(\mathcal{A}^{\prime})$ participate in the post-shattering random process and any event $E(v)$ or $E_{CC}(i)$ adjacent to such a vertex participates in the post-shattering as well; note that the event $E_{CC}(i)$ for a clique $A_{i}$ that has the closest vertex $w$ in $H^{\prime}$ in distance $2$ does not need to participate despite $w\in\operatorname{vbl}(E_{CC}(i))$ as the event is avoided regardless the color choice of $w$. Hence by Lemma 3.2, the connected components of $G_{\operatorname{dep}}$ have size at most $O(\Delta(F)^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}+c_{2}}\log n)=\operatorname{\text{{\rm poly}}}(\log n)$ with high probability.
@@
After retracting colors in $\operatorname{vbl}(\mathcal{A}^{\prime})$, none of the $E_{CC}(i)$ holds anymore and thus CC is maintained by the pre-shattering phase. During post-shattering, only vertices of $N_{H}^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))$ get colored. Hence, CC continues to hold for the $A_{i}$ that are not adjacent to some vertex in $N_{H}^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))$. If $A_{i}$ is adjacent to such a vertex, then CC is maintained by the post-shattering coloring (with a fresh budget) because $E_{CC}(i)$ would otherwise hold.
@@
$E_{CC}(i)$: CC is not maintained for some color w.r.t. $A_{i}$.
@@
We again apply Lemma 4.4 to upper bound the probability. Fix a clique $A_{i}$ and a color and form the sets consisting of external neighbors not in $\operatorname{Big}_{i}^{+}$ as before. Set $Q=\Delta^{1/10}$. The probability that a specific color is picked by neighbor is bounded above by $T/|L(v)|=1/\Delta^{7/20}\leqslant 1/(Q\cdot\Delta^{1/5})$. Due to the independence of picking candidate colors between different vertices the probability that a specific color is picked by $\ell$ vertices is at most $1/(Q\cdot\Delta^{1/5})^{\ell}$. No vertex lies in more than $2\Delta^{9/10}$ sets, because in that case it would belong to some $\operatorname{Big}_{i}^{+}$. Now we can apply Lemma 4.4 and get that for a fixed clique and color the probability that CC fails is at most $\exp(-\Delta^{1/40})$. With a union bound over all colors, $E_{CC}(i)$ holds w.p. at most $\Delta\cdot\exp(-\Delta^{1/40})$.
@@ 1. $E^{\prime}(v)$ for every uncolored vertex $v$, and
If $\Delta\geqslant(\log n)^{50}$, then by Lemmas 5.10 and 5.11, none of the bad events occur after running Algorithm 4 in pre-shattering, w.h.p.. Hence, $H$ is colored and CC is maintained w.h.p.. We henceforth assume that $\Delta\leqslant(\log n)^{50}$.
@@
Since the pre-shattering only colors vertices and that the slack does not decrease, Lemma 5.10 continues to apply to bound the probability of $E^{\prime}(v)$ in the post-shattering LLL. Likewise, Lemma 5.11 applies to bound the probability for $E_{CC}(i)$ in the post-shattering LLL. 5.9 bounds the dependency degree and hence, for sufficiently large $\Delta$, Algorithm 5 poses an LLL that we can solve with Theorem 3 in $\widetilde{O}(\log^{3}N)=\widetilde{O}(\log^{3}\log n)$ rounds; all other parts of the algorithm are executed in $O(1)$ rounds. At the end, all nodes of $H$ are colored as all events $E^{\prime}(v)$ are either avoided already after the pre-shattering phase or after the post-shattering phase; note that any uncolored node after the pre-shattering phase is contained in $\operatorname{vbl}(\mathcal{A}^{\prime})$. After retracting colors in $\operatorname{vbl}(\mathcal{A}^{\prime})$, none of the $E_{CC}(i)$ holds anymore and thus also CC is maintained by the pre-shattering phase. During post-shattering, only vertices of $\operatorname{vbl}(\mathcal{A}^{\prime})$ get colored. Hence, CC continues to hold for the $A_{i}$ that are not adjacent to some vertex in $\operatorname{vbl}(\mathcal{A}^{\prime})$. If $A_{i}$ is adjacent to such a vertex, then CC is maintained by the post-shattering coloring (with a fresh budget) as it is implied by avoiding $E_{CC}(i)$.
@@ 6 Coloring the Sparse Nodes
In this section, we show how to color the sparse nodes $S$ while maintaining $CC$. Recall that this is the first step of Algorithm 1.
@@ 6.1 Algorithm for Sparse Nodes
For sufficiently large $\Delta$, it holds that $\Delta^{3/4}/40>U_{1}\cdot\Delta^{0.22}$, thereby proving Part (b) of $\Pi$. As for Part (a), recall that a vertex in $A$ has at most $U_{1}=10^{8}\sqrt{\Delta}$ neighbors in $S$ (by Lemma 4.1(3)).
@@
2 Each vertex of $S^{\prime}$ joins $S_{2}$ independently w.p. $p$.
@@ 1. $E_{d}^{\prime}(v)$:
We use the following variant of the result in which a set $A$ of active nodes try colors from a custom set of $q$ colors, which we prove in Appendix C.
@@
There exists a universal constant $c_{0}$ for which the following holds. Let $\Delta$ and $q$ be positive integers with $q\geqslant\Delta/3$. Let $A\subseteq V(G)$ be a subset of nodes such that $F[A]$ has maximum degree at most $\Delta$, and CC a set of $q$ colors. Consider a node $v\in V$ with at least $\overline{m}$ non-edges in $F[N(v)\cap A]$ and such that $\overline{m}\geqslant c_{0}\cdot q$. If every vertex of $A$ samples a random color $\chi(v)\in C$ and retains it only if $\chi(v)\notin\chi(N(v)\cap A)$, then
@@
Suppose $H$ is a graph of maximum degree $\Delta$ and let $v$ be a vertex such that $H[N(v)]$ contains $\overline{m}$ anti-edges. Sample every vertex into a set $A$ with probability $p\geqslant 8/\Delta$. The number of anti-edges induced by $H[N(v)\cap A]$ is at least $p^{2}\overline{m}/2$ w.p. at least $1-\exp(-\Omega(p\overline{m}/\Delta))$.

### 01 Apr 2026 v2

@@ 5 Coloring With Much Slack
The goal of this section is to prove Lemma 4.5, that is, we aim to color an induced subgraph $H$ of $F-\bigcup_{i}A_{i}$ that is $\Pi$-ous while at the same time maintaining $C$ for all uncolored cliques. Recall that $H$ is $\Pi$-ous if each node has sufficient slack in $H$ compared to a given upper bound $U\geqslant\Delta^{1/4}$ on the external degree of uncolored cliques. The coloring process consists of two primary components. The first one is a random color trial for all uncolored vertices, applied for $O(\log\Delta)$ iterations, where each iteration lowers the uncolored degree of each vertex by a constant factor in expectation. The properties of the resulting partial coloring after a single color trial are summarized in Lemma 5.1 below. The second component summarized in Lemma 5.2 below is then a multi color trial that colors all remaining vertices of the graph instance to be colored.
@@
    $C$ is maintained for all uncolored cliques.
@@
Let $H$ be an uncolored induced subgraph in which every vertex has $\Delta^{9/20}$ slack and each uncolored vertex in each $A_{i}$ has at most $\Delta^{1/10}$ neighbors in $H$. There is a $\widetilde{O}(\log^{3}\log n)$-round $\mathsf{LOCAL}$ algorithm that, w.h.p., colors $H$ while maintaining $C$ for every uncolored clique $A_{i}$. When $\Delta\geqslant(\log n)^{50}$, the algorithm runs in $O(1)$ rounds.
@@
For sufficiently large constant $c$, this is at most $\Delta^{1/10}$. Now let $H^{\prime}$ be the subgraph of the initial graph $H$ that contains all the yet uncolored vertices after $T$ iterations of Lemma 5.1. Then the graph $H^{\prime}$ fulfills the preconditions of Lemma 5.2, which require that every vertex has $\Delta^{9/20}$ slack and each uncolored vertex in each $A_{i}$ has at most $\Delta^{1/10}$ neighbors in the remaining uncolored graph. The application of Lemma 5.2 to $H^{\prime}$ completes the coloring of $H$. The $C$ is preserved for all uncolored cliques as it is guaranteed by the outcome of Lemmas 5.1 and 5.2. The runtime consists of the $O(\log\Delta)$ iterations from Lemma 5.1 that take $\widetilde{O}(\log^{3}\log n)$ rounds each and $\widetilde{O}(\log^{3}\log n)$ rounds for the one application of Lemma 5.2. This completes the proof.
5.1 Iterated Random Color Trial with $C$ (Lemma 5.1)
The goal of this section is to prove Lemma 5.1. We present a $\operatorname{\text{{\rm poly}}}(\log\log n)$ algorithm that partially colors the graph $H$ while maintaining $C$ and additionally reduces the uncolored degree of vertices in $H$ and $\bigcup_{i}A_{i}$ below $\Delta^{1/10}$.
The analysis focuses on a single iteration of Random Color Trial (Rct). We show that in each iteration of the Rct procedure the uncolored degree of a vertex drops by a factor of (11/180)(1-1/180) compared to the previous iteration while maintaining $C$. We set up an LLL that, if solved correctly, yields a partial coloring guaranteeing this drop in the uncolored degree. The random process underlying our LLL is described in Algorithm 2; we call the color $\psi(v)$ the candidate color of $v$.
@@ 2 Each activated vertex $v$ independently samples a uniform available color $\psi(v)$
$E_{CC}(i)$: $C$ is not maintained for some color w.r.t. $A_{i}$.
@@
Consider any uncolored clique $A_{i}$ and any color $x$. The relevant collection of subsets is formed by the external neighborhoods of each vertex $v\in A_{i}$. By Lemma 4.1-(a), we have $|A_{i}|<\Delta$. We ignore vertices of $\mathrm{All}_{i}\cup\operatorname{Big}_{i}^{+}$ for their contribution is not accounted for in $C$. Hence, no relevant vertex of $H$ belongs to more than $2\Delta^{9/10}$ such subsets. Since $H^{\prime}$ is $\Pi$-ous, each vertex $v$ of $H^{\prime}$ picks color $x$ as candidate color w.p. at most $1/(\deg_{H^{\prime}}(v)+U\cdot\Delta^{0.22}))\leqslant 1/(Q\cdot\Delta^{1/5})$ for $Q=U$. As candidate color choices are independent for different vertices, the probability that $\ell$ vertices pick color $x$ is at most $1/(Q\cdot\Delta^{1/5})^{\ell}$. Therefore, by Lemma 4.4, $C$ fails for clique $A_{i}$ and color $x$ w.p. at most $\exp(-\Delta^{1/40})$. The lemma follows by the union bound over a total of at most $\Delta$ different colors.
@@
As detailed in $C$, there is a fresh budget of $CC(i,x)<\Delta^{\frac{37}{40}}$ for each color $x$, each clique $A_{i}$, and also for each coloring step of the algorithm. In particular, in Algorithms 3 and 3 of Algorithm 3 the events $E_{CC}(i)$ get a separate budget of $\Delta^{\frac{37}{40}}$ each.
@@
Step (S1) holds with $c_{1}=5$ due to 5.3 (and because $F$ includes the cliques): a vertex $v$ holds variables used only by events $E(u)$ with $u\in N_{H}^{\leqslant 2}(v)$ (which has diameter at most 4) or events $E_{CC}(i)$ with $A_{i}$ with $\operatorname{dist}_{G}(v,A_{i})\leqslant 2$ (which has diameter at most 5 in $F$). Step (S2) requires a bound $\Delta(F)^{c_{2}}$ for some constant $c_{2}$ on the number of events in which each variable appears and it holds with $c_{2}=2$ as each variable is only used by events with variables in its 22-hop neighborhood. Step (S3) requires that for each event $A\in\mathcal{A}$ we can determine $1(A\in\mathcal{A}^{\prime})$ by only evaluating the variables in $\operatorname{vbl}(A)$ which immediately holds by the definition of $\mathcal{A}^{\prime}$. Step (S4) requires an upper bound of $1/\Delta(F)^{c_{4}}$ on the probability for each bad event $A\in\mathcal{A}$ to be contained in $\mathcal{A}^{\prime}$ where $c_{1}$ is an arbitrary constant satisfying $c_{4}>3c_{1}(4c_{\mathcal{B}^{\prime}}+16)+c_{2}+c_{\mathcal{B}^{\prime}}+1$. By Lemmas 5.4 and 5.5 this probability is at most $\Delta(G)\cdot\exp(-\Delta(G)^{1/40})$. As $\Delta(G)=\Theta(\Delta(F))$ Step (S4) holds for $c_{4}$ (as chosen above) as long as $\Delta(G)$ is larger than a sufficiently large absolute constant $\Delta_{0}$. The set of events included in the LLL of Algorithm 3 are all included in $\mathcal{B}^{\prime}$ (with $c_{\mathcal{B}^{\prime}}$ defined above) as vertices with distance at most 22 to $\operatorname{vbl}(\mathcal{A}^{\prime})$ participate in the post-shattering random process and any event $E(v)$ or $E_{CC}(i)$ adjacent to such a vertex participates in the post-shattering as well; note that the event $E_{CC}(i)$ for a clique $A_{i}$ that has the closest vertex $w$ in $H^{\prime}$ in distance 22 does not need to participate despite $w\in\operatorname{vbl}(E_{CC}(i))$ as the event is avoided regardless the color choice of $w$. Hence by Lemma 3.2, the connected components of $G_{\operatorname{dep}}$ have size at most $O(\Delta(F)^{4c_{1}c_{\mathcal{B}^{\prime}}+4c_{1}+c_{2}}\log n)=\operatorname{\text{{\rm poly}}}(\log n)$ with high probability.
@@
After retracting colors in $\operatorname{vbl}(\mathcal{A}^{\prime})$, none of the $E_{CC}(i)$ holds anymore and thus $C$ is maintained by the pre-shattering phase. During post-shattering, only vertices of $N_{H}^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))$ get colored. Hence, $C$ continues to hold for the $A_{i}$ that are not adjacent to some vertex in $N_{H}^{\leqslant 2}(\operatorname{vbl}(\mathcal{A}^{\prime}))$. If $A_{i}$ is adjacent to such a vertex, then $C$ is maintained by the post-shattering coloring (with a fresh budget) because $E_{CC}(i)$ would otherwise hold.
@@
$E_{CC}(i)$: $C$ is not maintained for some color w.r.t. $A_{i}$.
@@
We again apply Lemma 4.4 to upper bound the probability. Fix a clique $A_{i}$ and a color and form the sets consisting of external neighbors not in $\operatorname{Big}_{i}^{+}$ as before. Set $Q=\Delta^{1/10}$. The probability that a specific color is picked by neighbor is bounded above by $T/|L(v)|=1/\Delta^{7/20}\leqslant 1/(Q\cdot\Delta^{1/5})$. Due to the independence of picking candidate colors between different vertices the probability that a specific color is picked by $\ell$ vertices is at most $1/(Q\cdot\Delta^{1/5})^{\ell}$. No vertex lies in more than $2\Delta^{9/10}$ sets, because in that case it would belong to some $\operatorname{Big}_{i}^{+}$. Now we can apply Lemma 4.4 and get that for a fixed clique and color the probability that $C$ fails is at most $\exp(-\Delta^{1/40})$. With a union bound over all colors, $E_{CC}(i)$ holds w.p. at most $\Delta\cdot\exp(-\Delta^{1/40})$.
@@ 1. $E^{\prime}(v)$ for every uncolored vertex $v$, and
If $\Delta\geqslant(\log n)^{50}$, then by Lemmas 5.10 and 5.11, none of the bad events occur after running Algorithm 4 in pre-shattering, w.h.p.. Hence, $H$ is colored and $C$ is maintained w.h.p.. We henceforth assume that $\Delta\leqslant(\log n)^{50}$.
@@
Since the pre-shattering only colors vertices and that the slack does not decrease, Lemma 5.10 continues to apply to bound the probability of $E^{\prime}(v)$ in the post-shattering LLL. Likewise, Lemma 5.11 applies to bound the probability for $E_{CC}(i)$ in the post-shattering LLL. 5.9 bounds the dependency degree and hence, for sufficiently large $\Delta$, Algorithm 5 poses an LLL that we can solve with Theorem 3 in $\widetilde{O}(\log^{3}N)=\widetilde{O}(\log^{3}\log n)$ rounds; all other parts of the algorithm are executed in $O(1)$ rounds. At the end, all nodes of $H$ are colored as all events $E^{\prime}(v)$ are either avoided already after the pre-shattering phase or after the post-shattering phase; note that any uncolored node after the pre-shattering phase is contained in $\operatorname{vbl}(\mathcal{A}^{\prime})$. After retracting colors in $\operatorname{vbl}(\mathcal{A}^{\prime})$, none of the $E_{CC}(i)$ holds anymore and thus also $C$ is maintained by the pre-shattering phase. During post-shattering, only vertices of $\operatorname{vbl}(\mathcal{A}^{\prime})$ get colored. Hence, $C$ continues to hold for the $A_{i}$ that are not adjacent to some vertex in $\operatorname{vbl}(\mathcal{A}^{\prime})$. If $A_{i}$ is adjacent to such a vertex, then $C$ is maintained by the post-shattering coloring (with a fresh budget) as it is implied by avoiding $E_{CC}(i)$.
@@ 6 Coloring the Sparse Nodes
In this section, we show how to color the sparse nodes $S$ while maintaining CCCC. Recall that this is the first step of Algorithm 1.
@@ 6.1 Algorithm for Sparse Nodes
For sufficiently large $\Delta$, it holds that $\Delta^{3/4}/40>U_{1}\cdot\Delta^{0.22}$, thereby proving Part (b) of $\Pi$. As for Part (a), recall that a vertex in AA has at most $U_{1}=10^{8}\sqrt{\Delta}$ neighbors in $S$ (by Lemma 4.1(3)).
@@
2 Each vertex of $S^{\prime}$ joins $S_{2}$ independently w.p. pp.
@@ 1. $E_{d}^{\prime}(v)$:
We use the following variant of the result in which a set AA of active nodes try colors from a custom set of $q$ colors, which we prove in Appendix C.
@@
There exists a universal constant $c_{0}$ for which the following holds. Let $\Delta$ and $q$ be positive integers with $q\geqslant\Delta/3$. Let $A\subseteq V(G)$ be a subset of nodes such that $F[A]$ has maximum degree at most $\Delta$, and $C$ a set of $q$ colors. Consider a node $v\in V$ with at least $\overline{m}$ non-edges in $F[N(v)\cap A]$ and such that $\overline{m}\geqslant c_{0}\cdot q$. If every vertex of AA samples a random color $\chi(v)\in C$ and retains it only if $\chi(v)\notin\chi(N(v)\cap A)$, then
@@
Suppose $H$ is a graph of maximum degree $\Delta$ and let $v$ be a vertex such that $H[N(v)]$ contains $\overline{m}$ anti-edges. Sample every vertex into a set AA with probability $p\geqslant 8/\Delta$. The number of anti-edges induced by $H[N(v)\cap A]$ is at least $p^{2}\overline{m}/2$ w.p. at least $1-\exp(-\Omega(p\overline{m}/\Delta))$.